the price of an item was rs 10 in 1990 its price has risen by rs 2 each year if this trend continues in which year the price will be 100
Answer by CubeyThePenguin(3113)

x = number of years passed since 1990

cost of item = 10 + 2x

10 + 2x = 100
x = 45

The price will be Rs. 100 in 1990 + 45 = 2035.